OpenStudy (anonymous):

convert 8.314 J/mol K to in^3 Pa/ lb mole R

OpenStudy (anonymous):

Break it up into discrete units... \[\frac{inches^3 \cdot Pa}{lb \cdot mol \cdot R}=\frac{inches^3}{lb} \cdot \frac{Pa}{mol \cdot R}\] \[\frac{J}{mol \cdot K}=\frac{m^3}{1}\cdot\frac{Pa}{mol \cdot K}\] \[J = \frac{kg \cdot m^2}{s^2} = N \cdot m = \rm Pa \cdot m^3={}\rm W \cdot s \] Based on what you wrote, I think you may have typed the question incorrectly.

OpenStudy (anonymous):

And yes, I'm well aware R = Rankine and K = Kelvin, measures for temperature from absolute zero.

OpenStudy (anonymous):

These are valid values for the gas constant: http://en.wikipedia.org/wiki/Gas_constant And thus the answer your are looking for can be found there.

OpenStudy (anonymous):

thank you :)